OpenStudy (anonymous):

Find the value of the constant K when:

OpenStudy (anonymous):

The coefficient of x^2 in the expansion of \[(k + \frac{ 1 }{ 3 } x)^5 = 30 \]

OpenStudy (vishweshshrimali5):

One way is to expand the expression but I won't advise it.

OpenStudy (vishweshshrimali5):

What I would suggest is using this instead: For the expansion of \((x+a)^n\), the rth term is given by: \[\large{\boxed{T_{r} = ^nC_{r-1}x^ra^{n-r}}}\] Now, first find out the term in which power of x is 2.

OpenStudy (anonymous):

This is what I got, not too sure about how to solve for r. Sometimes I get it, sometimes I don't. 5-2r = 2 -2r = -3 r = 3/2

OpenStudy (vishweshshrimali5):

Okay lets see: First of all I am going to simplify the expansion a little bit: \[\large{(k + \cfrac{1}{3}x)^5 = \cfrac{1}{3^5}(3k + x)^5}\] Now, since, the term you are asking for has power of x = 2 Thus, r = 2 Also, from the simplified expansion, a = 3k n = 5 So, the term would be: \[\large{T_2 = \cfrac{1}{3^5}\times ^5C_1 x^2 (3k)^{5-2}}\] Now, the coefficient part is clearly: \[\large{\cfrac{1}{3^5}\times^5C_1\times(3k)^3}\] (I am multiplying by 1/3^5 as it was multiplied by the whole expansion) But you are given that this is equal to 30. So, \[\large{\cfrac{1}{3^5}\times ~^5C_1\times (3k)^3 = 30}\] So, use this equation to find out k.

OpenStudy (anonymous):

Not sure if I'm right: \[\frac{ 1 }{3^5 } \times ^5C_{1} \times (3k)^3 = 30\] \[\frac{ 5 \times 27k^3 }{3^5 } = 30\] \[135k^3 = 7290\] \[k^3 = 54\] \[k = 3.78\]

OpenStudy (vishweshshrimali5):

You are correct (as far as I think)

OpenStudy (anonymous):

Given answer is 3 though, it could be the wrong answer - will check with my lecturer tomorrow. Thanks for the help! :)